If there are 'x' atoms in 5 grams of carbon, how many atoms are there in 5 grams of silicon?
1
Expert's answer
2017-01-23T06:48:38-0500
According to Avogadro’s law,
N = n · 6*1023
n = m/M
M (C) = 12 g/mol M (Si) = 28 g/mol
n (C) = 5/12 = 0.42 mol n (Si) = 5/28 = 0.18 mol
N (C) = 0.42 · 6*1023 = 2.52 *1023 atoms N (Si) = 0.18 · 6*1023 = 1.08 *1023 atoms